# Label-Guided In-Context Learning for Named Entity Recognition

### v1

@@ Abstract
In-context learning (ICL) enables large language models (LLMs) to perform new tasks using only a few demonstrations. In Named Entity Recognition (NER), demonstrations are typically selected based on semantic similarity to the test instance, ignoring training labels and resulting in suboptimal performance. We introduce Deer, a new method that leverages training labels through token-level statistics to improve ICL performance. Deer first enhances example selection with a label-guided, token-based retriever that prioritizes tokens most informative for entity recognition. It then prompts the LLM to revisit error-prone tokens, which are also identified using label statistics, and make targeted corrections. Evaluated on five NER datasets using four different LLMs, Deer consistently outperforms existing ICL methods and approaches the performance of supervised fine-tuning. Further analysis shows its effectiveness on both seen and unseen entities and its robustness in low-resource settings. [^1] Code and data are available at https://github.com/bflashcp3f/deer.
Label-Guided In-Context Learning for Named Entity Recognition
@@ 1 Introduction
Figure 1: Unlike existing in-context learning methods for NER that select demonstrations solely based on input embedding similarity, our proposed method, Deer, leverages token-level, label-related statistics to enhance demonstration retrieval in a training-free manner. It further uses this information to reflect on initial predictions and make refinements.
@@
Recent advances in large language models (LLMs; Brown et al., 2020; Touvron et al., 2023; Dubey et al., 2024; OpenAI et al., 2024) suggest an alternative: in-context learning (ICL; Dong et al., 2024), where models learn new tasks directly from examples provided as input, without parameter updates. As shown in Figure 1, a common ICL strategy for NER (Wang et al., 2023a; Lu et al., 2024) involves retrieving a fixed number of demonstrations using embeddings of input data, and then prompting the model to extract entities. While such methods show particular promise in low-resource scenarios (Lee et al., 2022), it does not scale as efficiently as supervised methods (Sainz et al., 2024). For example, Monajatipoor et al. (2024) report a 79.3 F1 score on the NCBI dataset using GPT-4 with 16 examples per query. In contrast, a fine-tuned BERT-base model achieves an 84.0 F1 score when given access to the same training data. This discrepancy suggests that the current ICL approach makes less effective use of the training data. For demonstration retrieval, similarity between the test example and training examples depends on sentence-level semantic similarity and does not use the label on the training instances. Task-agnostic sentence embeddings fail to capture token-level label details (e.g., entities), which are critical for NER. Similarly, the selected examples included in the prompt are the totality of label information provided to the model; the vast majority of annotations are ignored. As a result, the model’s predictions may miss known entities, invent spurious ones, or have inaccurate boundaries.
@@
Motivated by these findings, we propose Deer: Data statistics-grounded namEd Entity Recognition, an ICL method that leverages label statistics to support collective learning from the entire training set. Deer operates in two steps. First, it improves demonstration retrieval by aligning it with the token-level nature of NER. Rather than relying on task-agnostic sentence embeddings, it uses token-based retrieval guided by label statistics to prioritize tokens likely to belong to entities or their context, resulting in more relevant examples. Second, it introduces an error reflection step that identifies potentially misclassified tokens using the same token-level label statistics. These tokens are revisited with targeted span-level demonstrations, allowing the model to refine its predictions and improve overall performance.
@@
We evaluate Deer on five NER datasets spanning four domains: news, biomedicine, social media, and mixed web text, using four recent LLMs: Qwen2.5-7B, Llama3.3-70B, GPT-4o-mini, and GPT-4o. The results show that Deer consistently outperforms ICL baselines across datasets and models. With GPT-4o, Deer achieves performance competitive with supervised models on four out of five datasets, while the open-source Llama3.3-70B also demonstrates strong results. Further analysis confirms that Deer enhances example retrieval, improves recognition of both seen and unseen entities, and demonstrates robustness in low-resource settings.
@@ 2 Deer [Uncaptioned image]
We introduce Deer, an ICL approach designed to collectively leverage the full training labels while focusing on token-level details for NER. Deer comprises two steps, as illustrated in Figure 2: 1) label-guided retrieval, and 2) error reflection. The following sections detail the problem setup and our training data preparation process for Deer, followed by a description of each of its two steps.
2.1 Problem Setup and Training Data Preparation
@@
We support Deer through a pre-processing step: we categorize training input tokens into three distinct classes based on entity labels and compute the frequency and probability of each token within these classes (Step 0 in Figure 2). Entity tokens are those that belong to an identified entity. Context tokens are the tokens surrounding an entity, typically two tokens on each side, which provide identification cues and help define entity boundaries (Huang et al., 2015). Other tokens include all remaining tokens in the sentence. For example, consider the sentence in Figure 2, ‘‘The risk of cancer, ...’’, where “cancer” is an entity token, “risk” and “of” are context tokens, and “The” is an other token, as it falls out of the context window. Context tokens like “risk of” provide contextual clues for identifying Disease entities, even in cases where entities are unseen in the training data.
@@ 2.3 Error Reflection (ER)
Even with improved retrieval, ICL can only incorporate a small subset of training examples, leaving blind spots that limit the model’s ability to make well-informed predictions. For example, the model may miss entities present in the training set but absent from retrieved examples or make boundary errors due to insufficient guidance. To address this, we introduce an error reflection step that uses training statistics to identify error-prone tokens and provides targeted demonstrations to help the model revise its predictions. We focus on single-entity, boundary-related errors, which have been identified as a major source of NER mistakes (Ding et al., 2024; Lu et al., 2024). A hierarchical breakdown of NER errors is shown in Figure 4 in the appendix.
@@ 3.1 Dataset and Evaluation Metric
Table 1: Statistics of the five experimented datasets.
@@
Dataset Deerw/o ER Deer Deerw/o ER Deer BERT UniNER
NCBI 82.7 (\$2.1) 84.8 (\$3.0) 84.1 (\$6.5) 85.7 (\$7.3) 84.0 87.0
bc2gm 74.3 (\$2.9) 75.4 (\$3.7) 75.0 (\$9.4) 75.6 (\$9.9) 79.0 82.4
CoNLL03 93.5 (\$1.9) 93.6 (\$2.2) 93.9 (\$5.8) 94.0 (\$6.0) 90.4 93.3
OntoNotes 83.5 (\$2.5) 85.2 (\$3.3) 85.8 (\$7.5) 86.6 (\$8.2) 87.2 89.9
TweetNER7 63.6 (\$1.9) 64.3 (\$3.4) 64.9 (\$6.1) 65.3 (\$7.2) 60.9 65.7
Table 4: ICL results of Deer (using GPT-4o) with 8 and 32 demonstrations, compared to two fine-tuning baselines. Values in parentheses indicate associated API inference costs on the test set. Increasing the number of demonstrations improves performance, and error reflection continues to provide additional gains at larger scales. Notably, Deer with 8-shot error reflection performs competitively with the 32-shot setting without reflection, but at a significantly lower cost. Deer also matches or exceeds BERT-base on most datasets and narrows the gap with UniNER.
@@ 3.2 Baselines
We compare Deer against both ICL and fine-tuning baselines to demonstrate its improvements over standard ICL methods and its ability to close the performance gap with fine-tuning. All methods use the full training set for either fine-tuning or in-context demonstration selection.
@@
We include three ICL baselines: KATE (Liu et al., 2022), BM25 (Robertson et al., 2009), and BSR (Gupta et al., 2023). Kate serves as our primary ICL baseline, as it remains the dominant framework for ICL-based NER (Wang et al., 2023a; Berger et al., 2024; Monajatipoor et al., 2024). This method selects in-context examples through nearest neighbor search using sentence-level embeddings. BM25 treats both the test input and demonstrations as bags of words and measures their relevance via a TF-IDF–weighted recall of overlapping words. Recent IE work (Sun et al., 2024) has shown its effectiveness for demonstration retrieval. BSR employs BERTScore-Recall (Zhang et al., 2020) as a similarity metric, which measures token-level recall using token embeddings. While mainly used in semantic parsing, we adapt it here for NER. To ensure the competitiveness of the baselines, we evaluate seven embedding models commonly used in recent studies (see Appendix B). OpenAI’s text-embedding-3-small consistently outperforms open-source alternatives such as all-mpnet-base-v2, and is therefore used in all our experiments.
@@
We include two supervised baselines: BERT fine-tuning and UniversalNER-7B (Zhou et al., 2024). We fine-tune a BERT-base model (Devlin et al., 2019) for token-level classification with three random seeds and report the average performance. UniversalNER explores targeted distillation for NER, leveraging gpt-3.5-turbo to generate entity annotations for unlabeled text. The refined dataset is then used to train a student model, which is further fine-tuned on supervised NER datasets. This method achieves competitive results with state-of-the-art NER-specific LLMs, including InstructUIE (Wang et al., 2023b) and GoLLIE (Sainz et al., 2024). [^4] As noted in its GitHub repository, the authors have not released the code and prompts required to reproduce most of the supervised dataset results due to licensing restrictions. Therefore, we rely on the reported results from the original paper for comparison.
@@ 3.3 Implementation Details
We conduct experiments with four recent LLMs, covering both open-source and proprietary models: Qwen2.5-7B (Qwen et al., 2025), Llama3.3-70B (Dubey et al., 2024), GPT-4o-mini, and GPT-4o (OpenAI et al., 2024). [^5] In our preliminary experiments, we also evaluated Llama3.1-8B. However, it failed to generate outputs in the required format during ICL (see Appendix B.6 for more details). Thus, we did not pursue it further. We use the same prompt across four LLMs and set the decoding temperature to 0 to facilitate result reproducibility. For main ICL experiments, we retrieve eight demonstrations per query, sorted in ascending order of similarity to the test example. Appendix A contains detailed descriptions of the prompts and all hyperparameters.
@@ 4 Results
Table 2 presents the results of three ICL baselines and two variants of Deer, including the one without error reflection (Deerw/o ER), evaluated across five datasets and four LLMs, using eight retrieved demonstrations per query. Deerw/o ER consistently outperforms Kate in all settings, demonstrating the effectiveness of our label-guided, token-level demonstration retrieval. It also performs better than BM25 and BSR. Interestingly, BSR consistently outperforms Kate, suggesting that BERTScore-Recall, which emphasizes token-level coverage, may be more suitable than cosine similarity on sentence representations for token-centric tasks like NER. A promising future direction would be to integrate our label-guided retrieval with BERTScore-Recall. Adding error reflection leads to further gains across all settings, confirming its effectiveness. A breakdown of performance across the three reflection steps is provided in Table 7 in the appendix. Appendix C (Table 18) presents a detailed error analysis, highlighting Deer’s advantages in reducing under-predicted entities and boundary errors.
@@
To evaluate the scalability of Deer and its performance relative to fine-tuning, Table 4 reports results using GPT-4o with 8 and 32 demonstrations per query, along with two fine-tuning baselines. We include both performance and API inference costs, with and without error reflection. [^6] Inference costs are calculated using OpenAI’s official pricing: \$2.50 per million input tokens and \$10.00 per million output tokens. Increasing the number of demonstrations to 32 consistently improves performance across all datasets. Error reflection continues to provide added benefit even at this larger scale. Notably, Deer with error reflection and 8 demonstrations performs competitively with the 32-demo version without reflection, but at significantly lower cost. This highlights a potential cost-performance trade-off between scaling demonstrations and applying error reflection, which we leave for future work.
Compared to fine-tuning baselines, Deer matches or exceeds BERT-base on four of five datasets and narrows the performance gap with UniNER. These results suggest that reliance on task-specific NER models may decline over time, especially as general-purpose LLMs become more cost-efficient and can be reused across multiple tasks, while specialized models incur additional maintenance and deployment costs.
@@ 4.1 Ablation Studies & Analyses
To better understand the effectiveness of our retriever, we compare three retrieval methods across all five datasets using GPT-4o: sentence embeddings from Kate, our proposed token-based retriever, and an unweighted variant that excludes label-guided token statistics. All embeddings are generated using OpenAI’s text-embedding-3-small model. As shown in Table 5, our label-guided method consistently outperforms the other two. The unweighted variant performs similarly to or worse than sentence embeddings, underscoring the importance of incorporating label-aware, token-level weighting in retrieval. Further analysis of the two components in our retriever is provided in Appendix B.
Retriever Sentence Token (uwg.) Token (wg, ours)
@@
Table 5: Comparison of three demonstration retrievers: sentence embeddings (Kate), our label-guided token-based retriever, and an unweighted variant without label statistics. Embeddings are generated using OpenAI’s text-embedding-3-small. Our method consistently outperforms the others, highlighting the value of label-aware, token-level weighting.

### v2

# LLMs are Better Than You Think: Label-Guided In-Context Learning for Named Entity Recognition

@@ Abstract
In-context learning (ICL) enables large language models (LLMs) to perform new tasks using only a few demonstrations. However, in Named Entity Recognition (NER), existing ICL methods typically rely on task-agnostic semantic similarity for demonstration retrieval, which often yields less relevant examples and leads to inferior results. We introduce Deer, a training-free ICL approach that enables LLMs to make more informed entity predictions through the use of label-grounded statistics. Deer leverages token-level statistics from training labels to identify tokens most informative for entity recognition, enabling entity-focused demonstrations. It further uses these statistics to detect and refine error-prone tokens through a targeted reflection step. Evaluated on five NER datasets across four LLMs, Deer consistently outperforms existing ICL methods and achieves performance comparable to supervised fine-tuning. Further analyses demonstrate that Deer improves example retrieval, remains effective on both seen and unseen entities, and exhibits strong robustness in low-resource settings. [^1] Code and data are available at https://github.com/bflashcp3f/deer.
LLMs are Better Than You Think: Label-Guided In-Context Learning for Named Entity Recognition
@@ 1 Introduction
Figure 1: Existing ICL methods for NER typically select demonstrations based on task-agnostic input embeddings, which we argue yield suboptimal demonstrations for LLMs and consequently inferior performance. In contrast, our proposed method, Deer, leverages label-grounded statistics to identify entity-relevant tokens, enabling more task-focused demonstration selection and targeted error reflection.
@@
Recent advances in large language models (LLMs; Brown et al., 2020; Touvron et al., 2023; Dubey et al., 2024; OpenAI et al., 2024) suggest an alternative: in-context learning (ICL; Dong et al., 2024), where models learn new tasks directly from examples provided as input, without parameter updates. As shown in Figure 1, a common ICL strategy for NER (Wang et al., 2023a; Lu et al., 2024) involves retrieving a fixed number of demonstrations using embeddings of input data, and then prompting the model to extract entities. While such methods show particular promise in low-resource scenarios (Lee et al., 2022), it does not scale as efficiently as supervised methods (Sainz et al., 2024). For example, Monajatipoor et al. (2024) report a 79.3 F1 score on the NCBI dataset using GPT-4 with 16 examples per query. In contrast, a fine-tuned BERT-base model achieves an 84.0 F1 score when given access to the same training data.
In this paper, we argue that this performance gap primarily stems from inefficiencies in demonstration retrieval rather than from inherent limitations of LLMs. The prevailing sentence embedding–based retrievers are ill-suited for NER for two key reasons: (1) NER is a token-level task requiring fine-grained sensitivity to entity boundaries, whereas sentence embeddings operate at a coarse, sentence-level granularity; and (2) these embeddings are task-agnostic, disregarding label information from the training data, even though an ideal retriever should recognize entity-focused patterns and retrieve demonstrations containing similar entities to those in the query. As shown in Table 6, our sanity-check analysis confirms that sentence embedding–based retrievers frequently fail to retrieve demonstrations containing overlapping entities with the query, highlighting their inefficiency.
@@
Motivated by these findings, we propose Deer (Data Stat.-grounded namEd Entity Recognition), a method designed to help LLMs make more informed predictions during ICL in a training-free manner. The core idea of Deer is to leverage task-specific information embedded in label statistics. In the context of NER, this corresponds to identifying which tokens are more likely to appear within entity spans, near entity spans, or outside them. Based on such statistics from the entire training set, Deer incorporates them in two complementary stages. First, the retriever uses token-level statistics to guide demonstration selection, prioritizing examples containing tokens that are likely to correspond to entities or their contexts, resulting in more relevant demonstrations. Second, during an error reflection stage, Deer revisits potentially misclassified tokens (flagged by these statistics) with targeted span-level demonstrations, enabling the model to refine its predictions and achieve improved overall performance.
@@
We evaluate Deer on five NER datasets spanning four domains: news, biomedicine, social media, and mixed web text, using four LLMs: Qwen2.5-7B, Llama3.3-70B, GPT-4o-mini, and GPT-4o. The results show that Deer consistently outperforms ICL baselines across datasets and models. With GPT-4o, Deer achieves performance competitive with supervised models on four out of five datasets, while Llama3.3-70B also demonstrates strong results. Further analysis confirms that Deer enhances example retrieval, improves recognition of both seen and unseen entities, and demonstrates robustness in low-resource settings.
@@ 2 Deer [Uncaptioned image]
We introduce Deer, an ICL approach that enables LLMs to make more informed entity predictions through the use of label statistics. As illustrated in Figure 2, Deer consists of two main components: (1) label-guided retrieval and (2) error reflection. The following sections describe the problem setup and the construction of the label statistics in Deer, followed by explanations of each component.
2.1 Problem Setup
@@
2.2 Data Statistics Preparation
As motivated in §1, an effective demonstration retriever should capture entity-focused patterns and retrieve demonstrations that exhibit similar patterns to those in the query. To achieve this, we leverage label-grounded token statistics from the training data. Specifically, we categorize all training tokens into three classes based on their relation to entities and compute the frequency and probability of each token within these classes (Step 0 in Figure 2).
Entity tokens are those that belong to an identified entity. Context tokens are the tokens surrounding an entity, typically two tokens on each side, which provide identification cues and help define entity boundaries (Huang et al., 2015). For example, consider the sentence in Figure 2, ‘‘The risk of cancer, ...’’, where “cancer” is an entity token, and “risk” and “of” are context tokens. Context tokens like “risk of” provide contextual clues for identifying Disease entities, even in cases where entities are unseen in the training data. Finally, other tokens include all remaining tokens that fall outside the entity and its immediate context window (e.g., “The” in the example).
@@ 2.4 Error Reflection (ER)
Even with improved retrieval, ICL can only incorporate a small subset of training examples, leaving blind spots for LLM predictions. For example, the model may miss entities present in the training set but absent from retrieved examples or make boundary errors due to insufficient guidance. To address this, we introduce an error reflection step that uses prepared statistics to identify error-prone tokens and provides targeted demonstrations to help the model revise its predictions. We focus on single-entity, boundary-related errors, which have been identified as a major source of NER mistakes (Ding et al., 2024; Lu et al., 2024). A hierarchical breakdown of NER errors is shown in Figure 4.
@@ 3.1 Dataset and Evaluation Metric
Table 1: Statistics of the five experimented datasets. Each dataset contains a minimum of 5,000 annotated training examples, ensuring competitive performance of fine-tuned models.
@@
Dataset Deerw/o ER Deer Deerw/o ER Deer BERT UniNER7B InsUIE11B GoLLIE13B GoLLIE34B
NCBI 82.7 (\$2.1) 84.8 (\$3.0) 84.1 (\$6.5) 85.7 (\$7.3) 84.0 87.0 86.2 86.5 85.8
bc2gm 74.3 (\$2.9) 75.4 (\$3.7) 75.0 (\$9.4) 75.6 (\$9.9) 79.0 82.4 80.7 - -
CoNLL03 93.5 (\$1.9) 93.6 (\$2.2) 93.9 (\$5.8) 94.0 (\$6.0) 90.4 93.3 91.5 93.0 93.1
OntoNotes 83.5 (\$2.5) 85.2 (\$3.3) 85.8 (\$7.5) 86.6 (\$8.2) 87.2 89.9 88.7 84.0 84.6
TweetNER7 63.6 (\$1.9) 64.3 (\$3.4) 64.9 (\$6.1) 65.3 (\$7.2) 60.9 65.7 66.0 - -
Table 4: ICL results of Deer (using GPT-4o) with 8 and 32 demonstrations, compared to five fine-tuning baselines. Values in parentheses indicate associated API inference costs on the test set. Increasing the number of demonstrations improves performance, and error reflection continues to provide additional gains at larger scales. Notably, Deer with 8-shot error reflection performs competitively with the 32-shot setting without reflection, but at a significantly lower cost. Deer also matches or exceeds fine-tuning baselines on four out of five datasets.
@@ 3.2 Baselines
We compare Deer against both ICL and fine-tuning baselines to demonstrate its improvements over standard ICL methods and its ability to approach fine-tuning performance. All methods use the full training set for either fine-tuning or in-context demonstration selection.
@@
We cover three ICL baselines: KATE (Liu et al., 2022), BM25 (Robertson et al., 2009), and BSR (Gupta et al., 2023). Kate serves as our primary ICL baseline, as it remains the dominant framework for ICL-based NER (Wang et al., 2023a; Berger et al., 2024; Monajatipoor et al., 2024). This method selects in-context examples through nearest neighbor search using sentence-level embeddings. BM25 treats both the test input and demonstrations as bags of words and measures their relevance via a TF-IDF–weighted recall of overlapping words. Recent IE work (Sun et al., 2024) has shown its effectiveness for demonstration retrieval. BSR employs BERTScore-Recall (Zhang et al., 2020) as a similarity metric, which measures token-level recall using token embeddings. While mainly used in semantic parsing, we adapt it here for NER. To ensure the competitiveness of the baselines, we evaluate seven embedding models commonly used in recent studies (see Appendix B). OpenAI’s text-embedding-3-small consistently outperforms open-source alternatives such as all-mpnet-base-v2, and is therefore used in all our experiments. We do not include GPT-NER (Wang et al., 2023a), an entity-based retriever, due to its impracticality for the ICL setting. Nonetheless, we show in Table 14 that our method outperforms GPT-NER. Additional discussion on GPT-NER can be found in §5.
@@
We compare with five supervised models: BERT-base, UniversalNER-7B (Zhou et al., 2024), InstructUIE-11B (Wang et al., 2023b), and GoLLIE-13B/34B (Sainz et al., 2024). BERT-base (Devlin et al., 2019) is fine-tuned for token-level classification; we run three random seeds and report the averaged performance. UniversalNER-7B applies targeted distillation for NER by prompting gpt-3.5-turbo to annotate unlabeled text, training a student model on the refined data, and subsequently fine-tuning it on supervised NER datasets. [^4] As noted in its GitHub repository, the authors have not released the code or prompts necessary to reproduce most supervised results due to licensing restrictions. We therefore rely on the reported results in the original paper. InstructUIE-11B adopts unified instruction tuning, training FlanT5-11B (Chung et al., 2022) on 32 diverse IE datasets reformulated into a text-to-text format with expert-written instructions, yielding strong performance in both supervised and low-resource settings. Finally, GoLLIE-13B and GoLLIE-34B extend CodeLlama (Rozière et al., 2024) with human-authored guidelines, leading to substantial improvements in zero-shot IE while also maintaining competitive supervised results.
@@ 3.3 Implementation Details
We conduct experiments with four recent LLMs, covering both open-source and proprietary models: Qwen2.5-7B (Qwen et al., 2025), Llama3.3-70B (Dubey et al., 2024), GPT-4o-mini, and GPT-4o (OpenAI et al., 2024).555gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06 are used through OpenAI API. In our preliminary experiments, we also evaluated Llama3.1-8B. However, it failed to generate outputs in the required format during ICL (see Appendix B.6 for more details). Thus, we did not pursue it further. We use the same prompt across four LLMs and set the decoding temperature to 0 to facilitate result reproducibility. For main ICL experiments, we retrieve eight demonstrations per query, sorted in ascending order of similarity to the test example. Appendix A contains detailed descriptions of the prompts and all hyperparameters.
@@ 4 Results
Table 2 presents the results of three ICL baselines and two variants of Deer, including the one without error reflection (Deerw/o ER), evaluated across five datasets and four LLMs. Deerw/o ER consistently outperforms Kate in all settings, demonstrating the effectiveness of our label-guided, token-level demonstration retrieval. It also performs better than BM25 and BSR. Interestingly, BSR consistently outperforms Kate, suggesting that BERTScore-Recall, which emphasizes token-level coverage, may be more suitable than cosine similarity on sentence representations for token-centric tasks like NER. A promising future direction would be to integrate our label-guided retrieval with BERTScore-Recall. Adding error reflection leads to further gains across all settings, confirming its effectiveness. A breakdown of performance across the three reflection steps is provided in Table 8 in the appendix. Appendix C (Table 20) presents an error analysis, highlighting Deer’s advantages in reducing under-predicted entities and boundary errors.
@@
To evaluate the scalability of Deer and its performance relative to fine-tuning, Table 4 reports results using GPT-4o with 8 and 32 demonstrations per query, along with five fine-tuning baselines. We include both performance and API inference costs, with and without error reflection. [^6] Inference costs are calculated using OpenAI’s official pricing: \$2.50 per million input tokens and \$10.00 per million output tokens. Increasing the number of demonstrations to 32 consistently improves performance across all datasets. Error reflection continues to provide added benefit even at this larger scale. Notably, Deer with error reflection and 8 demonstrations performs competitively with the 32-demo version without reflection, but at significantly lower cost. This highlights a potential cost-performance trade-off between scaling demonstrations and applying error reflection, which we leave for future work.
Compared to five fine-tuning baselines, Deer achieves comparable or superior performance on four out of five datasets, e.g., surpassing GoLLIE-34B on CoNLL03 and OntoNotes. These results suggest that the reliance on task-specific NER models may diminish, as general-purpose LLMs become increasingly cost-effective and reusable across tasks, while specialized models incur additional maintenance and deployment overheads.
@@ 4.1 Ablation Studies & Analyses
To better understand the effectiveness of our retriever, we compare three retrieval methods across all five datasets using GPT-4o: sentence embeddings from Kate, our proposed token-based retriever, and an unweighted variant that excludes label-guided token statistics. All embeddings are generated using OpenAI’s text-embedding-3-small model. As shown in Table 5, our label-guided method consistently outperforms the other two. The unweighted variant performs similarly to or worse than sentence embeddings, underscoring the importance of incorporating label-grounded, token-level weighting in retrieval. Table 6 further provides a sanity check on demonstration retrieval quality: for entities that appear in both the training and test sets, our retriever includes a substantially higher proportion of these entities within the retrieved demonstrations compared to task-agnostic sentence embeddings, validating our argument about the inefficiency of generic retrievers. Further analysis of the two components in our retriever is provided in Appendix B.
Dataset Sentence Token (uwg.) Token (wg, ours)
@@
Table 5: Comparison of three demonstration retrievers: sentence embeddings (Kate), our label-guided token-based retriever, and an unweighted variant without label statistics. Embeddings are generated using OpenAI’s text-embedding-3-small. Our method consistently outperforms the others, highlighting the value of label-grounded, token-level weighting.
Dataset Sentence Token (wg, ours)
NCBI 85.0 93.3
bc2gm 72.5 83.1
CoNLL03 70.1 74.6
OntoNotes 62.7 80.6
TweetNER7 76.4 77.6
Table 6: Sanity check for demonstration retrievers. For entities appearing in both the training and test sets, we compare different retrievers based on the proportion of these entities contained within the retrieved eight demonstrations. Our retriever yields substantially more relevant entities than task-agnostic sentence embedding baselines, supporting our argument regarding the inefficiency of generic retrievers.